Let T : P2 → R2 be given by T(p) = p(1) +p(-1) p(2) Suppose that qe ker(T) is not 0. Find the roots of q.
Answer #1

Solution :

T : \mathbb{P} 2\rightarrow\mathbb{R}2 is given by, T(p) = [ p(1)+p(-1) p(2) ]t

q \in ker(T) and q \neq 0.

Thus,T(q) = [0 0]t and hence, [q(1) + q(-1) q(2)]t = [0 0]t

Thus, q(1)+q(-1) = 0 and q(2) = 0.

Since q \in \mathbb{P}2, hence there exist a,b,c in \mathbb{R} such that q(x) = ax2 + bx + c.  
Now, q(1)+q(-1) = 0 implies that a+b+c+a-b+c = 0 and hence, a = -c

Also, q(2) = 0 implies that 4a + 2b + c = 0. Using the relation a = -c, we obtain 3a = -2b.

Thus, we have 3a = -3c = -2b

Thus, q(x) = ax2 + bx + c = ax2 - 3ax/2 - a .

Hence, the roots of q are given by, ax2 - 3ax/2 - a = 0 which implies, 2x2 - 3x - 2 = 0. Thus, the roots of q are -1/2 and 2.